Solve the equation.
x(x-3)= 40

Answer by blwinbbbles(106)

First multiple out the equation:
x(x-3)=40 so: x^2 - 3x = 40
now set the equation zero:
x^2 - 3x -40 = 0 now you can factor the equation:
(x-8)(x+5) = 0
so either x-8 = 0 or x+5 = 0
so x is either 8 or -5...try them both back into the orginal equation
8(8-3)= 40
8(5)=40
40 = 40 yes
-5(-5-3) = 40
-5(-8)=40
40 = 40 yes
so the anwser is both x = 8 and x = -5